OpenStudy (anonymous):

Can you help me solve this: x^4-5x^2+2x+11=0

OpenStudy (anonymous):

Try using the Rational Root Theorem and Descartes' Rule of Signs.

OpenStudy (anonymous):

and what is that sorry! :)

OpenStudy (anonymous):

Rational Root Theorem: If we have a polynomial of the form: \[ P(x)=a_nx^n+a_{n-1}x^{n-1}+\cdots+a_0 \]Then, the only possible rational roots are those of the form, where \(p\) is any factor of \(a_0\) and \(q\) is a factor of \(a_n\): \[ x=\pm\frac{p}{q} \]So, in your case, our possible roots: \[ x=\{\pm\frac{1}{1}, \pm\frac{11}{1}\} \]Descartes' Rule of Signs is better by the net: http://www.purplemath.com/modules/drofsign.htm

OpenStudy (anonymous):

wow all this look very confusing can you help me solve my problem please =D @LolWolf

OpenStudy (anonymous):

All right, so, using synthetic division/plugging in numbers, we try every root possible. If \(x=1\): \[ (1)^4-5(1)^2+2(1)+11=0 \]Etc. And find that there are no real roots to the equation. And, we are done.

OpenStudy (anonymous):

really! we're done

OpenStudy (anonymous):

I mean to say that there are no *rational* solutions to the equation, just for future reference.

OpenStudy (anonymous):

ah okay i see